Can you please help. Find the slope and the y - intercept of the line. Write fractional answers in lowest terms.
2y + 8 x + 3 = 12 + 8 x.

Answer by mananth(16949)

2y + 8 x + 3 = 12 + 8 x.
2y+8x-8x=12-3
2y=9
/2
y=9/2.
This is a line parallel to the x axis and has no slope.
It intersects the y axis at 9/2.
...
m.ananth@hotmail.ca